OpenStudy (anonymous):

Rewrite this quadratic function f(x)=-2x^2 - 20x + 9 in standard form

OpenStudy (anonymous):

Multiply the whole quadratic by -1, Thus, f(x) = 2x^2 + 20x - 9

OpenStudy (anonymous):

and that is standard form?

OpenStudy (anonymous):

Yes. The x^2 term should always be positive. The standard form of a quadratic is ax^2 +/- bx +/- c.

OpenStudy (anonymous):

That is false. You cannot simply multiply the function by -1 as it becomes a different function than the original. The given function is in standard form. \(a\) is allowed to be negative.

OpenStudy (anonymous):

ok i was pulling my hair out trying to see what was going wrong here..........

OpenStudy (anonymous):

do you mean to put it in vertex form? 'cause as yakeyglee says, it already is in stadard form...

OpenStudy (anonymous):

it is telling me that it is wrong when i put it in as what it is already... it definitely says to put in standard form

OpenStudy (anonymous):

then maybe they are referring to vertex-form then. do you wanna begin?

OpenStudy (anonymous):

yes lets begin...

OpenStudy (anonymous):

btw, do you know what vertex form is?

OpenStudy (anonymous):

i am not sure how to do vertex

OpenStudy (anonymous):

i thought vertex was standard to be honest

OpenStudy (anonymous):

true... some people call vertex form as their standard form...

OpenStudy (anonymous):

you want to get your equation into looking like this: \[\large f(x) =a(x-h)^2+k\] where (h,k) is the vertex of your parabola.

OpenStudy (anonymous):

Hint: the axis of symmetry of a parabola is at \(-\frac{b}{2a}\), which is found by averaging the values of the zeroes output by the quadratic formula.

OpenStudy (anonymous):

so = a(x-20)+9

OpenStudy (anonymous):

f(x) = a(x - 20)^2 + 9 ?

OpenStudy (anonymous):

no... let's use the hint provided by yakeyglee... can you tell me the x-coordinate of the vertex. use your equation for the values of a and b.

OpenStudy (anonymous):

-20/-4

OpenStudy (anonymous):

good... so h=5 right?

OpenStudy (anonymous):

does that - before the equation cause the negative in the denominator to go away?

OpenStudy (anonymous):

the equation -b/2a

OpenStudy (anonymous):

yes... so -20/(2*(-2)) = 5

OpenStudy (anonymous):

this is the x-coordinate of your vertex.

OpenStudy (anonymous):

ok so i can plug h into that equation you gave me.

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

so far you have f(x)=-2(x-5)^2 + k

OpenStudy (anonymous):

all we gotta do is find k

OpenStudy (anonymous):

is k not given with -9?

OpenStudy (anonymous):

i mean + 9

OpenStudy (anonymous):

no. the k is the y-coordinate of the vertex. so find f(5) =

OpenStudy (anonymous):

-10 - 5^2 + k

OpenStudy (anonymous):

hang on man..... i was working with the wrong equation...

OpenStudy (anonymous):

i was working with -2x^2+20x-9 when it should have been -2x^2-20x-9

OpenStudy (anonymous):

why does that sign change from 9 to -9

OpenStudy (anonymous):

i got it wrong again.. sorry.... let's be clear... is this the function: f(x) = -2x^2 - 20x + 9

OpenStudy (anonymous):

that is the function. yes

OpenStudy (anonymous):

-b/2a = -(-20)/(2*(-2)) = -5 this is h. so h=-5 correct?

OpenStudy (anonymous):

-5 ok thats what i was wondering before.. yes.

OpenStudy (anonymous):

ok... we're good now... all we have to do is find k.

OpenStudy (anonymous):

to find k, you need to find f(h) = f(-5) =

OpenStudy (anonymous):

so F(h) = a(x+5)^2+k .

OpenStudy (anonymous):

what's k? find f(-5)

OpenStudy (anonymous):

plug in x=-5 into your equation...

OpenStudy (anonymous):

do you need help?

OpenStudy (anonymous):

ok i dont know what i did wrong.. i just got = a^2+k .. maybe im just not the best at evaluating

OpenStudy (anonymous):

f(-5) = -2*(-5)^2 - 20*(-5) + 9

OpenStudy (anonymous):

59

OpenStudy (anonymous):

ok... that's your k.. so a = -2 h = -5 k = 59 \[\large f(x)=-2(x-(-5))^2+59=-2(x+5)^2+59\]

OpenStudy (anonymous):

try that...

OpenStudy (anonymous):

that was it ! .. wow..

OpenStudy (anonymous):

so that equation -b/2a

OpenStudy (anonymous):

yes... that's to find the x coordinate of your vertex

OpenStudy (anonymous):

you can whatever that is to f and solve to get k... then solve from there

OpenStudy (anonymous):

set*

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

that is very helpful.. i can probably do these on my own now. thanks a lot

OpenStudy (anonymous):

sorry 'bout my messing up... could've finished this way faster if it wasn't for me... thanks for not getting pissed...

OpenStudy (anonymous):

yw... :)

OpenStudy (anonymous):

thank you yakey too, and yeah i just really wanted to learn how to do it, frustrated or not